OpenStudy (anonymous):

A round tablecloth hangs over the table an inch all around. If the table is 40 inches across, what is the area of the tablecloth? (1 point) 1384.74 square inches 1256 square inches 5024 square inches 131.88 square inches In cubic meters, how much greater is the volume of a cylinder with height 21m and a radius 12m than the volume of a square prism with the height 21 m and base 12 m on a side?

OpenStudy (anonymous):

maybe you could help me with this one too @saiberz

OpenStudy (anonymous):

If it hangs over an inch all around it means the radius of tablecloth is one inch higher than table radius Dt= 2Rt = 40 Rt = 20 Rtc=20+1 = 21 A=pi*r^2 Atc=pi*21^2=1385.44

OpenStudy (anonymous):

it is almost the same as 1384.74 ,the difference is beacuse how many digits of the number pi you use

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

yes i do wowo amazing thank you so much!

OpenStudy (anonymous):

you are welcome